A microscopic spherical dust particle of radius 2.1 μm and mass 9.5 μg is moving in outer space at a constant speed of 35 cm/s. A wave of light strikes it from the opposite direction of its motion and gets absorbed.

Assuming the particle decelerates uniformly to zero speed in one second, what was the maximum amplitude of the electric field of the radiation that was absorbed? Give your answer in volts per meter.
Answer #1

Partpt Force (F)=ma F=PA - a= acceleration ) m= mass ) (Psp nessure) A Area ma= PA =Par2 - P=ma tre = P = m uz (2 4 =) pt = m